OpenStudy (anonymous):

Round each number to the nearest hundredth. What is the best estimate for the sum of 12.547 + 83.919? A. 95.46 B. 95.47 C. 96.46 D. 96.47

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

Uh 12.547+83.919 =12.547+83.919 =96.466

OpenStudy (anonymous):

Round each number to the nearest tenth. What is the best estimate for the difference of 264.72 128.49? A. 137 B. 136.3 C. 136.2 D. 130

OpenStudy (anonymous):

@Thehulk49

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

Here 264.72-128.49=136.23